# InvBench: Can LLMs Accelerate Program Verification with Invariant Synthesis?

### v1

@@ 5.1 Results of LLMs
[…]
Qwen2.5-72B 4.4% 0.9% 1.20× 1.00×
gpt-oss-120b 8.8% 5.3% 1.10× 1.01×
claude-sonnet-4 15.0% 8.8% 1.06× 1.01×
Qwen3-Coder-480B 14.2% 8.0% 1.09× 1.01×
claude-opus-4.1 18.6% 8.8% 1.23× 1.02×
gpt-5 37.2% 26.5% 1.24× 1.06×
o3 39.8% 28.3% 1.37× 1.09×
[…]
gpt-5 11.5% 0% 1.00× 1.00×
Qwen2.5-72B 12.4% 0% 1.00× 1.00×
o3 29.2% 0% 1.00× 1.00×
Qwen3-Coder-480B 15.9% 0% 1.00× 1.00×
claude-sonnet-4 13.3% 0.9% 3.35× 1.01×
claude-opus-4.1 14.2% 0.9% 2.96× 1.02×
gpt-oss-120b 11.5% 0.9% 29.57× 1.03×
Table 3: InvBench-Hard: Results on the InvBench dataset (hard split) across models.
We report the performance of different LLMs on the easy and hard splits of InvBench in Table˜2 and Table˜3, respectively. On the easy split shown in Table˜2, o3 achieves the strongest results, with a 1.37× speedup over UAutomizer on 28.3% of problems and an overall 1.09× average speedup. These results indicate that while state-of-the-art LLMs generally struggle to synthesize correct or sufficiently strong invariants to accelerate program verification, the strongest model o3 can still yield non-trivial improvements on a meaningful fraction of problems.
Table˜3 reports results on the hard split of InvBench. In this setting, LLMs show only negligible improvement over UAutomizer. The only noteworthy exception is gpt-oss-120b, which produces an invariant for one problem that delivers a 29.57× speedup. However, overall, such speedup cases are exceedingly rare.
There are three main takeaways from the results. First, generating strong invariants that yield performance speedups is substantially more difficult than merely producing correct invariants, as reflected in the large gap between the percentage of correct invariants and the percentage of speedups. Second, model capability is a key factor, as demonstrated by the sharp differences in speedups reported in Table˜2. Third, LLMs remain far from fully addressing the task of invariant synthesis, leaving considerable room for future progress, as shown in Table˜3.
5.2 Results of Verifiers on InvBench
We compare the state-of-the-art non-LLM-based tool UAutomizer (Schüssele et al., 2024) with other LLM-based verifiers on both the easy split and the hard split of InvBench.
On the easy split of InvBench, where UAutomizer solves all 113 instances within 30 seconds, it clearly surpasses all LLM-based verifiers. LaM4Inv (Wu et al., 2024a) and Loopy (Kamath et al., 2023) fail to solve any instance within this time limit, while LEMUR (Wu et al., 2024b) solves only 55 instances. These results indicate that none of the LLM-based verifiers provide any advantage over UAutomizer on the easy split.
Table˜4 presents the comparison on the hard split of InvBench. UAutomizer still delivers the best performance across almost all timeouts. UAutomizer consistently and significantly outperforms LaM4Inv and Loopy. In particular, LaM4Inv and Loopy solve fewer than half as many instances as UAutomizer across all timeouts. Notably, within 10 seconds, LEMUR solves more problems than UAutomizer, suggesting that it can accelerate some of the instances. Nevertheless, UAutomizer still outperforms LEMUR with longer timeouts, indicating that although LEMUR can offer early solving advantages on certain challenging instances, its overall capability is still limited compared to UAutomizer.
Given the consistently poor performance of LaM4Inv, which often either fails or times out, we conducted a manual investigation. Our analysis shows that LaM4Inv fails on some examples when applied to external datasets beyond those used in its customized evaluation. However, the preprocessing steps or manual annotations required by the tool are not documented. The official repository does not provide preprocessing scripts, and our attempt to contact the authors has not received a response. For a direct comparison between UAutomizer and LaM4Inv, we also refer readers to Section˜5.3, where we evaluate UAutomizer on the dataset released by LaM4Inv.
5.3 Results of UAutomizer on Prior Datasets
As shown in Table 5, UAutomizer, the state-of-the-art non-LLM-based verifier, consistently solves more instances than LaM4Inv (Wu et al., 2024a) and Loopy (Kamath et al., 2023), two representative LLM-based tools, on their respective customized datasets. This highlights that prior work omitted an important baseline comparison against UAutomizer.
LEMUR (Wu et al., 2024b) is the only tool that surpasses UAutomizer. However, this is largely explained by its dataset construction: LEMUR evaluates only on problems that UAutomizer cannot solve within 600 seconds. To assess whether this advantage generalizes beyond its own curated benchmark, we conduct an additional analysis presented in Section˜6.
While LaM4Inv and Loopy consistently underperform UAutomizer in terms of the total number of solved instances across time budgets, they nevertheless offer complementary benefits, which we discuss in Section˜A.2.
We further investigate the distributional differences between datasets. At the 600-second timeout, the performance of prior LLM-based verifiers approaches that of UAutomizer, with LaM4Inv even matching it exactly. In contrast, on InvBench (see Section˜5.2), all LLM-based verifiers perform poorly at 600 seconds, suggesting a substantial shift in distribution compared to the customized datasets used in prior work. Our analysis confirms this: programs in InvBench are significantly longer, averaging 62 lines of code in the hard split, compared to only 22 for LaM4Inv, 23 for LEMUR, and 27 for Loopy. Manual inspection shows that InvBench contains features such as multiple loops, functions, arrays, and pointers, which are largely absent from prior datasets. This makes InvBench a more challenging and realistic benchmark that better distinguishes solver performance.
We also note that different LLM-based verification frameworks are built on different base solvers: LEMUR is built on UAutomizer, Loopy on Frama-C (Cuoq et al., 2012), and LaM4Inv on ESBMC (Gadelha et al., 2018). Given the strength of UAutomizer, we believe future work should place greater emphasis on developing LLM-based verifiers atop state-of-the-art solvers such as UAutomizer and ensure that comparisons against it are not omitted.
Verifier Solved Instances under Different Timeouts
10s 100s 300s 600s
LaM4Inv (Wu et al., 2024a) 0 0 0 0
Loopy (Kamath et al., 2023) 0 2 18 50
LEMUR (Wu et al., 2024b) 32 71 73 73
UAutomizer 0 73 95 113
Table 4: Comparison of solved instances on InvBench-Hard by different verifiers under varying timeout budgets.
@@ 5.3 Results of UAutomizer on Prior Datasets
Table 5: Comparison of prior LLM-based verifiers and UAutomizer on their own customized dataset under different timeout budgets.
5.4 Performance Gains from Invariants Generated by UAutomizer
To quantify the potential speedup achievable when providing correct and strong invariants to solvers, we conduct a study using UAutomizer. We extract the invariants identified by UAutomizer and then measure the resulting speedup when they are supplied to the verifier. On a random sample of 100 problems from our training set, we observe an overall average speedup of 1.86×. This indicates that the invariants discovered by UAutomizer are sufficiently strong to accelerate verification, and we leverage them for fine-tuning LLMs in Section˜5.5.
5.5 Effectiveness of Fine-Tuning
For each synthetically generated problem in the training set, we run UAutomizer to produce invariants, yielding a training dataset of 3589 problems with associated invariants. We then perform supervised fine-tuning using LoRA for 3 epochs.
[…]
Qwen2.5-72B (base) 4.4% 0.9% 1.20× 1.00×
Qwen2.5-72B (fine-tuned) 32.7% 26.5% 1.13× 1.03×
Qwen3-Coder-480B (base) 14.2% 8.0% 1.09× 1.01×
Qwen3-Coder-480B (fine-tuned) 40.7% 29.2% 1.29× 1.08×
Table 6: Improvement of supervised fine-tuning on InvBench-Easy.
Table˜6 reports results on the easy split of InvBench. Fine-tuning leads to substantial gains in both invariant correctness and runtime speedup. Qwen3-Coder-480B shows the most pronounced improvements: the proportion of correct invariants increases from 14.2% in the base model to 40.7%, and the percentage of instances with speedups rises from 8% to 29.2%. For Qwen2.5-72B, the conditional speedup decreases after fine-tuning, but this is because the baseline conditional speedup was driven by a single case, whereas fine-tuning yields a larger number of cases with speedups. We also note that on the hard split, neither of the fine-tuned models shows improvement. These results suggest that more training efforts are required to achieve gains on harder instances.
5.6 Effectiveness of Best-of-N Sampling
[…]
claude-opus-4.1 1 18.6% 8.8% 1.23× 1.02×
claude-opus-4.1 16 37.2% 23.0% 1.15× 1.03×
claude-sonnet-4 1 15.0% 8.8% 1.06× 1.01×
claude-sonnet-4 16 36.3% 22.1% 1.20× 1.04×
o3 1 39.8% 28.3% 1.37× 1.09×
o3 16 50.4% 35.4% 1.39× 1.12×
Table 7: Results of Best-of-N sampling (N = 16) on InvBench-Easy.
We also evaluated whether repeated sampling improves LLM performance in invariant synthesis. For each problem, we generated 16 samples at a temperature of 0.7, removed duplicate invariants, and verified all candidates in parallel. We then select the sample with the shortest verification time and kill all other samples. Table 7 summarizes the results on InvBench’s easy split, demonstrating that repeated sampling leads to consistent gains. For example, o3 improves from 39.8% to 50.4% correct invariants, while claude-opus-4.1 nearly doubles its accuracy from 18.6% to 37.2%.
On the hard split, for claude-opus-4.1, with 16 samples, the percentage of instances with correct invariants increases from 14.2% to 15.9%, the percentage of instances with speedup from 0.9% to 2.7%, and the overall average speedup from 1.02× to 1.03×. Similarly, for claude-sonnet-4, the percentage of instances with correct invariants increases from 13.3% to 20.4%, the percentage of instances with speedup from 0.9% to 2.7%, conditional average speedup from 3.35× to 6.12×, and the overall average speedup from 1.01× to 1.07×. These improvements suggest that best-of-N sampling is a promising technique for LLMs to improve invariant synthesis performance.
6 Discussion
Generalizability of LEMUR.
LEMUR (Wu et al., 2024b) reports the best results on its customized dataset, as it targets instances that UAutomizer cannot solve within 600 seconds. To test whether this advantage generalizes, we sampled 50 unsolved instances and found that LEMUR solved 12 within the same timeout of 600 seconds. This suggests that LEMUR’s gains are not solely due to benchmark design but reflect a genuine advantage on harder problems.
Inference Overhead of Models.
Our evaluation includes LLM serving time as a realistic measure of end-to-end performance. Since the goal is to accelerate verification, inference overhead must be considered alongside solver runtime. Future research may explore how to balance the quality of generated invariants with the inference cost of producing them.
7 Conclusion
This work introduced InvBench, a principled framework for evaluating the capabilities of LLMs in invariant synthesis. Our approach employs a verifier-based decision procedure with a formal soundness guarantee and assesses not only correctness but also the speedups that invariants contribute to program verification. Using a benchmark of 226 instances, we conducted a comparison across state-of-the-art LLMs, existing LLM-based verifiers, and the traditional solver UAutomizer. The results show that although LLM-based verifiers represent a promising direction, they do not yet offer significant advantages over non-LLM-based approaches. Model capability proves to be a critical factor, and our benchmark remains an open challenge for current LLMs. At the same time, we demonstrated that supervised fine-tuning and Best-of-N sampling can improve model performance in accelerating verification.
LLM Usage
LLMs are the subject of this study. We additionally used them for polishing the writing.
Appendix A Appendix
A.1 Proof
Theorem (Decision Soundness).
Proof.
For the $\mathbf{T}$ case, the final rule must be DEC-PROP with $d=\mathbf{T}$. The premises yield $V(P,\varnothing,q)=\mathbf{T}$ and $V(P,\{q\},p^{\star})=\mathbf{T}$. By verifier soundness, $P\models q$ and $P\models_{\{q\}}p^{\star}$. Since $q$ holds on all executions of $P$, introducing the assumption $\{q\}$ does not remove executions relevant to $p^{\star}$; thus $P\models p^{\star}$.
For the $\mathbf{F}$ case, the final rule must be DEC-FALSE. Its premise $V(P,\{q\},p^{\star})=\mathbf{F}$ implies $P\not\models_{\{q\}}p^{\star}$ by soundness. Assumptions restrict behaviors; hence, a violation under assumptions entails a violation without them, yielding $P\not\models p^{\star}$. ∎
A.2 Complementary Results of LLM-based Solvers
As shown in Table˜A1, all three LLM-based tools solve instances that remain unsolved by UAutomizer within the 600-second budget. LaM4Inv and Loopy each contribute additional solved cases, and LEMUR is able to handle 19 problems that UAutomizer cannot solve at all, underscoring the complementary strengths of LLM-based approaches.
Verifier Solved Instances $\Delta$ Solved
LaM4Inv 299 13
Loopy 403 40
LEMUR 19 19
Table A1: LLM-based verifiers complement UAutomizer by solving problems beyond its reach. We report the total numbers of solved problems, and $\Delta$ Solved” is the number of instances uniquely solved that are unsolved by UAutomizer.

### v2

@@ 5.1 Results of LLMs
Table˜2 presents the performance of different LLMs when using Quokka. The results are sorted by the number of instances achieving at least $2.0\times$ speedup (# Fast2.0), which ranges from 18 instances for Qwen2.5-7B to 51 instances for gpt-5.2. Overall, the table demonstrates substantial variation in both correctness and speedup performance across different models, with larger and more recent models generally achieving better results.
Effectiveness of Quokka.
Our approach demonstrates the speedup capabilities across all evaluated models. Even smaller models like Qwen2.5-7B achieve speedups on 34 instances at the $1.2\times$ threshold, while the best-performing model gpt-5.2 accelerates 81 instances at this level and achieves $2.0\times$ or greater speedup on 51 instances. These results demonstrate that Quokka can effectively leverage LLMs to generate invariants that meaningfully accelerate verification. The consistent speedup gains across different model families, from 7B parameter models to frontier models, and the progressive improvement with model scale underscore the effectiveness of our approach.
Correctness vs. Performance Gap.
A striking observation is the substantial gap between generating correct invariants and achieving meaningful speedups. For example, gpt-5.2 produces 710 correct invariants but achieves $2.0\times$ speedup on only 51 instances, while claude-opus-4.5 generates 689 correct invariants yet delivers $2.0\times$ speedup on merely 34 instances. This gap highlights that correctness alone is insufficient: invariants must also be strong enough to meaningfully accelerate verification, requiring not only logical correctness but also strategic insight into which invariants will most effectively guide the solver.
Model Inference Time.
We include model inference time in our evaluation to provide a fair and realistic assessment of end-to-end performance. As documented in Appendix A.4, smaller models such as Llama-3.1-8B and Qwen2.5-7B exhibit negligible generation times (0.3–0.4 seconds), while most models complete inference within 3 seconds. The o3 model has the highest overhead at 8.1 seconds on average. Importantly, unlike LOOPY (Kamath et al., 2023), which invokes the LLM up to 15 times per program and thus spends a substantial portion of its time budget on API calls, our approach requires only a single LLM invocation per program. Consequently, LLM inference accounts for only a small fraction of total runtime. Furthermore, our evaluation dataset contains sufficiently challenging programs (with solving times up to 600 seconds) such that solver time naturally dominates inference overhead, making the comparison meaningful even when inference latency is included.
Prompting-Based Methods.
We investigate the impact of different prompting strategies on invariant generation quality using gpt-5.2. As detailed in Appendix A.6, we evaluate 0-shot, 2-shot, 4-shot, and chain-of-thought (CoT) prompting. Few-shot prompting does not consistently improve performance, while CoT prompting yields the weakest results, likely due to increased generation time without proportional quality gains. These findings indicate that simple zero-shot prompting is most effective for invariant synthesis.
5.2 Comparison with Prior LLM-based Verifiers
Method # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
LaM4Inv 1 1 1 1 1
LOOPY 8 7 5 3 1
LEMUR-original 19 16 13 12 10
LEMUR-gpt-5.2 39 30 28 25 22
Quokka-gpt-5.2 81 65 60 56 51
Table 3: Comparison of speedup performance across different methods. We report the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over UAutomizer.
Table˜3 compares the speedup performance of Quokka against prior LLM-based verifiers across different speedup thresholds, while Figure˜2 shows the number of solved instances over time for each method.
Quokka Outperforms Prior LLM-based Verifiers.
Quokka with gpt-5.2 achieves state-of-the-art results across all speedup metrics. Among prior work, LEMUR (Wu et al., 2024b) represents the strongest baseline. To ensure a fair comparison, we enhance LEMUR by replacing its original model (gpt-4) with gpt-5.2, creating LEMUR-gpt-5.2. Even with this enhancement, Quokka-gpt-5.2 significantly outperforms LEMUR-gpt-5.2: at the $1.2\times$ speedup threshold, Quokka-gpt-5.2 accelerates 81 instances versus 39 for LEMUR-gpt-5.2 and 19 for LEMUR-original. At $2.0\times$ speedup, Quokka-gpt-5.2 achieves 51 instances compared to 22 for LEMUR-gpt-5.2 and 10 for LEMUR-original. LaM4Inv (Wu et al., 2024a) and LOOPY (Kamath et al., 2023) demonstrate minimal speedup capabilities, achieving at most 1 and 8 instances respectively at the $1.2\times$ level. These results demonstrate that Quokka establishes a new state-of-the-art for LLM-based verification acceleration.
Differences in Base Solvers.
Different LLM-based verification frameworks are built on different base solvers: Quokka and LEMUR use UAutomizer, LOOPY uses Frama-C (Cuoq et al., 2012), and LaM4Inv uses ESBMC (Gadelha et al., 2018). The choice of base solver significantly impacts the performance, as stronger solvers provide a better foundation for speedup. Since UAutomizer represents the current state-of-the-art solver, we believe future work should prioritize developing LLM-based verifiers atop state-of-the-art solvers such as UAutomizer. For a direct comparison between UAutomizer and prior LLM-based solvers, we also refer readers to Appendix A.3, where we evaluate UAutomizer on the dataset released by each of the prior works.
Algorithmic Simplicity and Parallelism.
Prior work predominantly adopts complex, sequential, and iterative processes with highly customized algorithms to handle LLM-generated invariants. In contrast, Quokka employs a simple, verifier-based, and parallel approach that directly leverages the base solver’s capabilities.
LaM4Inv uses a “query-filter-reassemble” strategy that filters and combines model-generated predicates through a complex algorithm, rather than using them directly. LOOPY employs a highly-customized Houdini algorithm with iterative fixing to filter candidates. LEMUR is closest to our approach, also directly using the base solver to verify the correctness of invariants, but follows a purely sequential decision process with backtracking. In contrast, Quokka directly uses LLM-generated invariants without decomposition or reassembly, and exploits parallelism by issuing the verification queries in parallel. This design reduces implementation complexity while enabling more efficient hardware utilization, contributing to the superior speedup performance demonstrated in our experiments.
Figure 2: Number of instances solved by different methods over varying timeout budgets.
Model # Incorrect Invariants # Goal Timeout # Invariant Timeout # Both Timeout
gpt-5.1 78 96 245 110
claude-sonnet-4 110 160 209 145
gpt-5.2 56 139 230 121
Table 4: Failure mode breakdown for the top three models on Quokka. Incorrect Invariants: the invariant is refuted by the solver. Goal Timeout: the invariant is verified, but proving the assertion under it times out. Invariant Timeout: verifying the invariant itself times out. Both Timeout: both verification queries time out.
Model # Correct Invariants # Fast1.0 # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
Qwen2.5-7B (base) 419 116 34 26 23 20 18
Qwen2.5-7B (finetuned) 431 141 39 27 26 24 20
Table 5: Performance comparison of base and fine-tuned Qwen2.5-7B. Fine-tuning leads to modest improvements in both correctness and speedup metrics.
5.3 Failure Mode Analysis
To understand why many invariants fail to accelerate verification in Quokka, we analyze failure modes for the top three models. We categorize failures into four types: 1) Incorrect Invariants: the invariant is refuted by the solver; 2) Goal Timeout: the invariant is verified, but proving the assertion under it times out; 3) Invariant Timeout: verifying the invariant itself times out; 4) Both Timeout: both verification queries time out. The results in Table˜4 reveal that while some failures stem from incorrect invariants, the dominant issue is timeout-related. This suggests that current models lack an understanding of what makes verification queries tractable for symbolic solvers. Future work should explore training reward models or other mechanisms to help models internalize solver difficulty, enabling them to propose invariants that genuinely simplify verification by decomposing hard problems into easier subproblems rather than inadvertently generating verification conditions that are as difficult or harder to solve than the original proof goal.
Figure 3: Effect of Best-of-N sampling on the number of instances under different Fastp settings.
5.4 Results of Fine-Tuning and Best-of-N Sampling
Modest Improvement from Supervised Fine-Tuning.
We perform supervised fine-tuning using LoRA (Hu et al., 2022) on Qwen2.5-7B for 3 epochs. As shown in Table˜5, fine-tuning yields modest improvements: correct invariants increase from 419 to 431, and instances achieving meaningful speedups improve across all thresholds. For example, Fast1.0 improves from 116 to 141, and Fast1.2 improves from 34 to 39. These results suggest that domain-specific fine-tuning can improve both correctness and speedup, though gains remain moderate.
Benfits of Best-of-N Sampling.
We evaluate Best-of-N sampling by generating multiple invariant candidates from gpt-5.2 at temperature 0.7 and selecting the one yielding the fastest verification time. As shown in Figure˜3, increasing N from 1 to 4 consistently improves speedup across all Fastp thresholds. However, performance plateaus beyond N=8 due to hardware constraints: our 20-core CPU must verify 2N queries in parallel, and when N exceeds 10, resource contention may negate sampling benefits. Memory limits can further constrain scalability. While Best-of-N sampling is effective, its benefits are hardware-bounded, though modest values of N yield meaningful improvements.
6 Conclusion
This work introduced Quokka, a simple yet effective framework for LLM-based invariant synthesis that accelerates program verification. Our approach employs a principled verification procedure that decomposes the original proof goal into two independent subproblems and launches verifier queries in parallel, providing sound evaluation while achieving state-of-the-art speedup results. Using a benchmark of 866 instances, we evaluated 9 state-of-the-art LLMs across multiple model families and compared against prior LLM-based verifiers. The results demonstrate that Quokka consistently outperforms all existing LLM-based verifiers, achieving speedups of at least 1.2× on 81 instances compared to 39 instances for the previous best approach. We further demonstrated that supervised fine-tuning and Best-of-N sampling can yield measurable improvements in accelerating verification. Our work establishes a new state-of-the-art for LLM-based program verification and provides a principled foundation for future research in this direction.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Appendix
A.1 Prompt Template Used for Training Dataset Generation
Table˜A1 shows the prompt template used for synthesizing training programs from seed programs.
You will be shown 3 example C programs. Please gain inspiration from the following programs to create a new high-quality C program. Do not simply copy from any of them. Requirements for the generated program: 1. The program MUST contain non-trivial loops (for or while). 2. The program MUST contain assertions. 3. The program MUST be compilable, self-contained, and reasonably complex (not trivial or overly short). 4. Only output the new C program. Example snippets: Program 1: {SEED_PROGRAM_1}
Program 2:
{SEED_PROGRAM_2}
Program 3:
{SEED_PROGRAM_3}
Output format: The generated program must be wrapped strictly in the following format: ‘‘‘c <NEW_C_PROGRAM>
Table A1: Prompt template for synthetic data generation.
A.2 An Example from the Fine-Tuning Dataset
Original Program (Input to UAutomizer)
1int main() {
2 int N = __VERIFIER_nondet_int();
3 assume_abort_if_not(N >= 1 && N < 100);
4 int x = 0, y = 0;
5 for (int i = 0; i < N; i++) {
6 if (i % 2 == 0) {
7 x += i;
8 } else {
9 y += i;
10 }
11 }
12 int diff = x - y;
13 __VERIFIER_assert(diff <= N);
14 return 0;
15}
Loop Invariant Generated by UAutomizer:
1Line Number: 6
2Predicate:
3(
4(i + 1) % 2 == 0 &&
5x < 2 + i + y &&
6x < 2 + y + 2 * i &&
7x < N + y + 1 &&
81 <= N
9)
10||
11(
12i % 2 == 0 &&
13x < 2 + i + y &&
14x < 2 + y &&
151 <= N
16)
Figure A1: An example from the fine-tuning dataset: program and its loop invariant generated by UAutomizer.
Figure˜A1 shows an example from the fine-tuning dataset with the program to the UAutomizer and the generated loop invariant.
The loop invariant holds at Line 6 (the beginning of the loop). It is a disjunction of two clauses, and can be written as $I\equiv P\lor Q=\big((i+1)\bmod 2=0\land x<2+i+y\land x<2+y+2i\land x<N+y+1\land 1\leq N\big)\lor\big(i\bmod 2=0\land x<2+i+y\land x<2+y\land 1\leq N\big)$.
By inspecting the loop, we can derive an exact relationship between $x$, $y$, and $i$ at the beginning of each iteration. Since $x$ accumulates all even numbers less than $i$ and $y$ accumulates all odd numbers less than $i$, we obtain:
$\text{if $i$ is even:}\quad x-y=-\frac{i}{2},\qquad\text{if $i$ is odd:}\quad x-y=\frac{i-1}{2}.$
Using this relationship, it is straightforward to verify that the invariant $I$ produced by UAutomizer is correct. Moreover, $I$ is strong enough to prove the final assertion $x-y\leq N$.
To show that the assertion holds at loop termination, we can check the following verification condition
$I\;\wedge\;(i=N)\;\Rightarrow\;(x-y\leq N),$
where $i=N$ denotes the loop’s exit condition.
Case 1: $N$ is odd.
Instantiating the invariant with $i=N$ activates the $P$ disjunct of $I$, from which we obtain
xy<N+1.x-y<N+1.
Since xyx-y is an integer, this directly implies $x-y\leq N$.
Case 2: $N$ is even.
In this case, the $Q$ disjunct applies. From the clause $x<2+y$ contained in $Q$, we derive
xy<2.x-y<2.
Because xyx-y is an integer and $N\geq 1$ and even (hence $N\geq 2$), we conclude
$x-y\leq 1\leq N,$
establishing the desired post-condition.
Thus, the invariant $I$ indeed suffices to prove the final assertion.
A.3 Results of UAutomizer on Prior Datasets
@@
Table A2: Comparison of prior LLM-based verifiers and UAutomizer on their own custom dataset under different timeout budgets.
As shown in Table A2, UAutomizer, the state-of-the-art non-LLM-based verifier, consistently solves more instances than LaM4Inv (Wu et al., 2024a) and Loopy (Kamath et al., 2023), two representative LLM-based tools, on their respective custom datasets. This demonstrates the importance of base solver choice.
LEMUR (Wu et al., 2024b) is the only tool that surpasses UAutomizer on its own benchmark. This outcome is largely explained by LEMUR’s dataset construction methodology: it specifically curates problems that UAutomizer cannot solve within 600 seconds. To assess whether LEMUR’s advantage generalizes beyond this carefully selected benchmark, we conduct an additional analysis. We randomly selected 50 instances that UAutomizer failed to solve within 600 seconds and evaluated LEMUR’s performance on them. LEMUR successfully solved 12 of these 50 instances within the same 600-second timeout, demonstrating that its gains are not merely an artifact of benchmark design but reflect genuine capability on challenging problems. Nevertheless, for our work, we focus on problems that UAutomizer can solve within the time budget, as our primary objective is to quantify speedup.
A.4 Inference Time of Different Models
Table˜A3 presents the average inference time per instance for different LLMs when generating invariants. As expected, smaller open-source models such as Llama-3.1-8B and Qwen2.5-7B exhibit the fastest generation times, averaging around 0.3–0.4 seconds per instance. Mid-sized models like Qwen2.5-72B and the GPT-5 series require approximately 0.6–1.4 seconds. The Claude models take longer at around 1.6–2.7 seconds per instance. The o3 model has the highest inference overhead, averaging over 8 seconds per instance. These results highlight the trade-off between model capability and inference cost, which must be considered when evaluating end-to-end verification performance.
Model Avg Generation Time (s)
Llama-3.1-8B 0.300
Qwen2.5-7B 0.355
Qwen2.5-72B 0.578
gpt-5.2 1.076
gpt-5.1 1.373
claude-opus-4.1 1.626
claude-opus-4.5 2.538
claude-sonnet-4 2.746
o3 8.075
Table A3: Average inference time per instance for different LLMs when generating invariants.
Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
gpt-5.1 399 10 9 9 9 9
gpt-5.2 455 17 11 10 9 9
Table A4: Performance of different LLMs when using ESBMC as the base verifier. We report the number of instances with verified-correct invariants and the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over ESBMC.
A.5 Results for ESBMC as the Base Solver
To demonstrate the generality of Quokka’s methodology, we conduct additional experiments using ESBMC (Gadelha et al., 2018) as the base solver instead of UAutomizer. As shown in Table˜A4, Quokka successfully accelerates ESBMC as well, with both gpt-5.1 and gpt-5.2 generating hundreds of correct invariants and achieving speedups on multiple instances. For example, gpt-5.2 produces 455 correct invariants and achieves speedups of at least $1.2\times$ on 17 instances. These results confirm that Quokka is not limited to a specific solver and can be applied to accelerate different verification tools.
A.6 Results for Prompting-Based Methods
Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
gpt-5.2 (CoT) 639 41 32 26 22 20
gpt-5.2 (4-shot) 756 72 62 53 47 43
gpt-5.2 (2-shot) 764 83 67 58 53 49
gpt-5.2 710 81 65 60 56 51
Table A5: Comparison of different prompting strategies for gpt-5.2.
We investigate the impact of different prompting strategies on invariant generation quality using gpt-5.2. As shown in Table˜A5, we compare chain-of-thought (CoT) prompting, few-shot prompting with 2 and 4 examples, and our default zero-shot approach. The results show that few-shot prompting does not necessarily improve performance and may introduce noise or distract the model from the task at hand. Chain-of-thought prompting yields the weakest results across all metrics, with only 639 correct invariants and 41 instances achieving $1.2\times$ speedup. This suggests that explicit reasoning steps may increase generation time without proportional benefits to invariant quality, thereby reducing overall speedup. These findings indicate that prompting-based strategy can hardly improve the task of invariant synthesis.